OpenStudy (anonymous):

The ticket office at Orchestra Center estimates that if it charges x dollars for box seats for a concert, it will sell 50 - x box seats. The function S = 50x - x2 gives the estimated income from the sale of box seats. Graph the function, and use the graph to find the price for box seats that will give the greatest income.

OpenStudy (anonymous):

You may want to know what the maximum income S(x) is: dS/dx = 50 - 2x = 0; x = 25 S(25) = 25 * 25 = 625 The maximum income is $625. The ticket price for that is $25.

OpenStudy (anonymous):

@jenniferlynn92 Questions, comments?